# Token Interdependency Parsing (Tipping) - Fast and Accurate Log Parsing

### v1

@@ 3.4 Interdependency Graph
Before going into the interdependency graph, it’s essential to define dependency in the context of this paper. Token dependency, inspired by recommender systems’ support concept [31] and conditional probability, is defined as the ratio of log messages in which both tokens appear together to the number of log messages that contain the dependent one. More in-depth, the dependency of the token $A$ to $B$ is the co-occurrence(s) of $A$ and $B$ together divided by the total occurrences of $A$. To formalize this, if $S_{A}$ represents the set of messages containing token $A$ while $S_{B}$ denotes those with token $B$, the dependency of token $A$ on token $B$ can be mathematically expressed based on these occurrences as follows:
@@
On the other hand, the interdependency graph is a directed graph constructed for each log message individually, where nodes (vertices) are tokens and edges (arcs) are the dependence of the source nodes’ tokens on destination nodes’ tokens.
@@ 3.5 Strongly Dependent Token Group
A strongly dependent token group is a subset of tokens within a message characterized by their mutual dependence on one another. Conversely, tokens outside this subset remain independent from the group, meaning either they do not depend on the tokens within the group or the group tokens do not depend on them. It is worth mentioning that there is at least one token group in a message, while the maximum number is equal to the number of unique tokens in that message.
The intuition behind grouping tokens based on their dependency is that when token $A$ depends on token $B$, and $A$ is identified as part of the template, inevitably, $B$ is also part of the template as token $A$ is unlikely to occur without token $B$. Therefore, in our context, tokens within a token group are expected to appear together.
A threshold hyper-parameter, denoted as $\theta$, is required to compute strongly dependent token groups (reasons are going to be discussed later). The process begins by constructing a token interdependency graph from the message. Then, edges with weights below the $\theta$ are then disregarded. Next, strongly dependent token groups are identified using strongly connected component algorithms, such as the Tarjan [32] or Kosaraju-Sharir [33] algorithm.
There are two reasons for disregarding edges below the threshold. First, it helps eliminate low dependencies often caused by noise. Second, since any two tokens within a single message are mentioned together at least once, a baseline of dependency always exists among all tokens in a message, leading to a fully connected interdependency graph. On the other hand, strongly connected components algorithms focus only on the presence of edges, not their weights, which eventually results in a single strongly dependent token group.
Algorithm 1 outlines all the steps required to compute the strongly dependent token groups.
@@ 3.6 Anchor Tokens Selection
After identifying strongly dependent token groups, the largest one is selected as the anchor group, and its members are designated as the anchor tokens.
@@ 3.7 Anchor-token-based Clustering
After identifying anchor tokens, messages are clustered into distinct groups according to their anchor tokens. Then, event IDs could be obtained from clusters by computing cluster hash or its index. Should the desired outcome be solely the event IDs, the algorithm concludes at this juncture. However, if there is a request for templates or parameter masks in addition to event IDs, the algorithm proceeds to produce the next three optional steps.
@@ 3.8 Cluster Common Tokens
Common tokens within each cluster are those present across all messages in the cluster, effectively representing the intersection of tokens in all messages. Subsequently, these common tokens are later leveraged to generate event templates and parameter masks, offering a refined approach to parsing and analysis.
@@ 3.9 Event Templates
Unlike previous methods, which produced a single template for each unique event, our method produces a set of templates for each event to handle the dynamicity occurring in log messages.For instance, Table I illuminates an example from the Proxifier dataset. In the example, all messages are representations of the same event. However, due to optional parameters within the log message structure[1], templates differ from each other.
The process of generating templates starts with a cluster of log messages and their common tokens. Initially, it involves processing these messages to create preliminary templates, where common tokens are retained, and the remaining elements are replaced with placeholders. To streamline the sequence, consecutive placeholders are merged into a single one. Subsequently, to make the template representable as a string, placeholders are replaced by a specific parameter symbol, such as “<*>”, to indicate variable parts. The final step aggregates these templates into a set, producing the templates for that cluster.
@@ 3.10 Parameter Masks
After identifying the common tokens across a cluster of messages, we initialize the parameter mask with an empty string for each message. Then, the parameter mask is populated by iterating over the message tokens, appending $n$ zeros if the token is present in the common tokens or $n$ ones otherwise, where $n$ is the number of characters in each token. Repeating this for all messages generates parameter masks for the entire cluster.
@@ 4 Experiments
Given that our algorithm relies on parallel computing, we implemented our algorithm in Rust due to its fearless concurrency, exceptional performance, and seamless compatibility with other languages. However, since Python is the dominant language in data science and AI space, we also released a Python version of our package (the core is still written in Rust), which will be utilized in the upcoming experiment. Tipping Python package can be installed from pypi [^1] https://pypi.org/project/tipping, while the Rust package is available on crates.io [^2] https://crates.io/crates/tipping-rs.
We performed four experiments, see Sections 4.1 to 4.4, to answer the following research questions:
  • RQ1: How accurate is Tipping in log parsing? Experiments in Sections 4.1 and 4.2
  • RQ2: How fast is Tipping in log parsing? Experiments in Sections 4.2 and 4.3
  • RQ3: What is the impact of Tipping as a parser on the down-the-stream task? Experiment in Section 4.4
4.1 LogPai’s Benchmark
TABLE II: A detailed comparison of the Tipping model’s performance against other leading state-of-the-art models in LogPai’s benchmark using the Group Accuracy as the evaluation metric.
Dataset Tipping Brain NuLog Drain Spell MoLFI Logram ULP
Android 0.94 0.96 0.83 0.91 0.92 0.63 0.79 0.84
Apache 1.00 1.00 1.00 1.00 1.00 1.00 0.31 1.00
BGL 0.98 0.99 0.98 0.96 0.79 0.96 0.59 0.93
HDFS 1.00 1.00 1.00 1.00 1.00 1.00 0.93 1.00
HPC 0.99 0.94 0.95 0.89 0.65 0.81 0.91 0.95
Hadoop 0.95 0.95 0.95 0.78 0.95 0.45 0.99
HealthApp 1.00 1.00 0.88 0.78 0.64 0.32 0.27 0.90
Linux 0.97 1.00 0.69 0.60 0.29 0.36 0.36
Mac 0.91 0.94 0.82 0.79 0.76 0.62 0.57 0.81
OpenSSH 0.81 1.00 0.79 0.55 0.54 0.61 0.43
OpenStack 1.00 1.00 0.97 0.73 0.76 0.21 0.33 0.49
Proxifier 1.00 1.00 0.53 0.53 0.00 0.50 0.02
Spark 1.00 1.00 1.00 0.92 0.90 0.42 0.28 0.92
Thunderbird 0.98 0.97 0.96 0.84 0.64 0.55 0.68
Windows 1.00 1.00 0.99 1.00 0.99 0.41 0.69 0.41
Zookeeper 0.99 0.99 0.97 0.96 0.84 0.72 0.99
TABLE III: A detailed comparison of the Tipping model’s performance against other leading state-of-the-art models in LogPAI’s benchmark using $\mathbf{F_{1}}$ score as the evaluation metric.
Dataset Tipping Brain NuLog Drain Spell MoLFI Logram ULP
Android 1.00 1.00 0.97 1.00 0.99 0.85 0.98 0.97
Apache 1.00 1.00 1.00 1.00 1.00 1.00 0.64 1.00
BGL 1.00 1.00 1.00 1.00 0.96 1.00 0.96 1.00
HDFS 1.00 1.00 1.00 1.00 1.00 1.00 0.99 1.00
HPC 1.00 1.00 0.99 0.99 0.99 0.98 0.99 0.99
Hadoop 1.00 1.00 1.00 0.92 1.00 0.78 1.00
HealthApp 1.00 1.00 1.00 0.92 0.89 0.78 0.59 0.99
Linux 1.00 1.00 0.99 0.94 0.72 0.54 0.48
Mac 0.99 1.00 0.75 0.98 0.96 0.93 0.89 0.98
OpenSSH 1.00 1.00 1.00 0.92 1.00 0.98 0.94
OpenStack 1.00 1.00 1.00 0.99 0.99 0.73 0.74 0.83
Proxifier 1.00 1.00 0.78 0.83 0.74 0.75 0.74
Spark 1.00 1.00 1.00 0.99 0.99 0.51 0.48 0.99
Thunderbird 1.00 1.00 1.00 0.99 1.00 0.99 1.00
Windows 1.00 1.00 1.00 1.00 1.00 0.91 0.91 0.99
Zookeeper 1.00 1.00 1.00 1.00 1.00 0.92 1.00
@@ 4.1.1 Experiment depiction
LogPai’s benchmark [24], published in 2019, has become a canonical accuracy evaluation benchmark for log parsers since its introduction. The benchmark comprises 2,000 manually labeled samples across 16 public open datasets. It primarily uses group accuracy and the $F_{1}$ score as its main metrics. Additionally, it includes time measurement during parsing as a performance metric. However, it is observed that performance might not be fully representative of real-world scenarios due to the limited number of samples in each dataset[1]. It’s important to note that although parsers are evaluated using larger datasets in the original paper, to the best of our knowledge, these larger datasets have not been made available in the GitHub repository.
LogPai permits the tuning of hyper-parameters specific to each dataset. However, we allocated no more than five minutes to tune the parameters for each dataset, as Tipping can reach maximum accuracy provided with the appropriate settings. Moreover, we were cautious not to overcomplicate the tuning process, aiming to maintain as much generalizability as possible in our approach.
It’s worth mentioning that since both group accuracy and $F_{1}$ score metrics are hardware-agnostic, we did not need to rerun the other state-of-the-art algorithms with our benchmark machine. Instead, we borrowed the results from the public GitHub repository [^3] https://github.com/logpai/logparser.
In favor of transparency, we made our implementation of Tipping in LogPai’s benchmark, which is a fork of the main repository, publicly available [^4] https://github.com/M3SOulu/logparser. Furthermore, we intend to merge this with the original repository after publishing the paper.
@@ 4.1.2 Experiment results
In Table II, Tipping demonstrates strong performance across various datasets, achieving perfect scores (1.00) in multiple cases such as Apache, HDFS, HealthApp, Proxifier, Spark, and Windows. While it shows slightly lower performance in others, like Android, BGL, HPC, Hadoop, Linux, Mac, OpenSSH, OpenStack, Thunderbird, and Zookeeper, Tipping’s scores are consistently high, mostly above 0.90, with the lowest score being 0.81 in the OpenSSH dataset. The results illustrate Tipping’s robustness to a diverse range of log datasets. Its performance is competitive compared to other state-of-the-art parsers, often matching or outperforming them in almost all datasets, underscoring its effectiveness as a log parsing tool.
Table III showcases the $F_{1}$ scores achieved by log parsers in the LogPai’s benchmark, where Tipping and Brain both scored consistently at 1.00 across almost every dataset, showcasing their top-tier accuracy. NuLog, however, showed variability in performance in the available datasets as its results were not available for all datasets. Other parsers like Spell, MoLFI, Logram, and ULP showed more variability in their scores, with Tipping outperforming them, highlighting its effectiveness as a log parsing tool compared to other state-of-the-art approaches.
@@ 4.2.1 Experiment depiction
While LogPai has gained widespread popularity within the field of log parsing as a benchmark for evaluating the accuracy of log parsing methods, it lacks significant features. The benchmark’s reliance on group accuracy as the defining metric and containment of insufficient quantity of samples to accurately reflect real-world performance scenarios have prompted researchers to seek alternatives. In response to these limitations, the recently published LogPM benchmark[1] has been introduced, addressing some of the criticisms of LogPai. LogPM distinguishes itself by emphasizing the detection of parameters at the character level. More in-depth, drawing inspiration from the field of computer vision, specifically semantic segmentation, LogPM introduces a new output for log parsers called the ”Parameter Mask.” This novel approach is accompanied by a new metric, the Parameter Mask Agreement (PMA), designed to compare the produced parsers’ parameter masks against the ground truth. Furthermore, LogPM endeavors to better replicate real-world prerequisites for performance evaluation by offering significantly larger datasets than those found in LogPai.
In favor of transparency, we merged our implementation of Tipping in LogPM benchmark into the main publicly available repository [^5] https://github.com/M3SOulu/LogPMBenchmark.
@@ 4.2.2 Experiment results
TABLE IV: An in-depth comparison of the Tipping’s performance with other leading state-of-the-art parsers in the LogPM benchmark, utilizing Parsing Accuracy as the evaluation metric.
@@
In Table IV, Tipping demonstrates a consistently higher parsing accuracy performance across a variety of datasets compared to its counterparts: Brain, Drain, Spell, and Lenma, as indicated by its average accuracy of 0.71. While Brain and Lenma show strong performances in specific datasets, they lack the overall consistency of Tipping. Particularly, Brain’s strengths in certain areas like HPC and Hadoop are counterbalanced by weaker performances in others, and Lenma, despite its high marks in HPC and Zookeeper, does not maintain this level of accuracy across all datasets. On the other hand, Drain and Spell exhibit lower overall accuracies. For instance, Drain excels in the SSH dataset but falls short in broader applications, indicating a niche area of strength. Spell’s performance is even less consistent, further emphasizing Tipping’s advantage in delivering higher accuracy across diverse datasets. Overall, Tipping’s performance underscores its reliability and capacity to adapt to varied data types, making it a preferred choice for comprehensive LogPm parsing tasks where data variability is a significant factor.
TABLE V: An in-depth comparison of the Tipping’s performance with other leading state-of-the-art parsers in the LogPM benchmark, utilizing Parameter Mask Agreement as the evaluation metric.
@@
In Table V, presenting LogPM’s parameter mask agreement scores, Tipping emerges as a leading method with an average agreement score of 0.87, showcasing its superior performance across various datasets. This performance is indicative of Tipping’s effectiveness in accurately identifying and matching parameter masks, a critical aspect of modern log parsing[34]. Comparatively, Brain and Lenma each hold an average agreement score of 0.80, indicating a commendable performance but still trailing behind Tipping. Both methods show strengths in certain areas, such as Lenma’s impressive accuracy in Hadoop and Brain’s exceptional performance in SSH, illustrating their potential in specific contexts. However, Tipping’s across-the-board excellence demonstrates a more universally applicable solution to parameter mask agreement challenges. Drain and Spell, with average scores of 0.72 and 0.71, respectively, lag further behind, particularly struggling in datasets like Android and Apache, where Tipping excels. Overall, Tipping’s performance is one level above its competition.
TABLE VI: An in-depth comparison of the Tipping’s performance with other state-of-the-art parsers in the LogPM benchmark’s speed test. The numbers are normalized for each dataset by dividing them by the lowest value, making the comparison between different datasets size-agnostic.
Dataset Tipping Brain Drain Spell Lenma
HPC 2.25 3.00 1.00 21.45 22.38
Zookeeper 2.17 4.37 1.00 13.16 30.98
Android 1.00 1.47 7.26 16.68 4435.41
Apache 2.98 6.69 1.00 13.68 126.45
Hadoop 1.00 1.22 3.52 21.06 9.03
HDFS 1.00 2.15 29.85 22.61 11.42
Linux 1.80 1.67 1.00 23.03 22.59
Openstack 2.79 3.74 1.00 43.43 44.24
Proxifier 2.58 3.35 1.00 6.75 23.67
SSH 2.83 6.88 1.00 17.77 42.89
Avg. 1.00 1.69 2.33 9.78 233.73
In Table VI, Tipping demonstrates remarkable efficiency, with an average speed score that is the fastest performance among the parsers, illustrating its exceptional speed in processing log data across various datasets. This performance is especially noteworthy in datasets like Android, Hadoop, and HDFS, where Tipping matches or outpaces its competitors by a significant margin. Meanwhile, Brain and Drain offer competitive speeds in certain datasets, but they do not consistently match Tipping’s rapid processing times across the board. Spell and Lenma, on the other hand, show significantly slower parsing speeds in comparison, with Lenma’s performance in the Android dataset highlighting a stark difference in efficiency. Lenma’s average speed score is drastically higher due to its outlier performance in handling the Android dataset, indicating a potential inefficiency in processing log data with a large number of templates. Tipping’s uniform speed across diverse datasets underscores its efficiency and reliability, making it capable of swiftly handling large volumes of log data and ideal for various environments.
4.3 Computational Efficiency with LogLead

### v2

@@ 3.4 Interdependency Graph
Before defining the interdependency graph, it’s essential to define dependency in this context. Token dependency, inspired by the concept of support in recommender systems[lu2015recommender] and conditional probability, is defined as the ratio of log messages in which both tokens appear together to the number of log messages that contain the dependent one. In other words, the dependency of the token $A$ on $B$ is the co-occurrence(s) of $A$ and $B$ together divided by the total occurrences of $A$. To formalize this, if $S_{A}$ represents the set of messages containing token $A$ while $S_{B}$ denotes those with token $B$, the dependency of token $A$ on token $B$ can be mathematically expressed as:
@@
Accordingly, the interdependency graph is a directed graph constructed for each log message individually, where nodes (vertices) are tokens and edges (arcs) are the dependence of the source nodes’ tokens on destination nodes’ tokens.
@@ 3.5 Strongly Dependent Token Group
A strongly dependent token group is a subset of tokens within a message characterized by their direct or indirect mutual dependence on each other. Conversely, tokens outside this subset either do not depend on the tokens within the group or the group tokens do not depend on them. It is worth mentioning that there is at least one token group in a message, while the maximum number is equal to the number of unique tokens in that message.
The intuition behind dependency grouping is that when token $A$ depends on token $B$, and $A$ is identified as part of the template, inevitably, $B$ is also part of the template, as token $A$ is unlikely to occur without token $B$. Therefore, tokens within the same token group are expected to appear together.
A threshold hyper-parameter, denoted as $\theta$, is required to compute strongly dependent token groups (reasons are going to be discussed in the next paragraph). The process begins by constructing a token interdependency graph from the message. Then, edges with weights below the $\theta$ are disregarded. Then, strongly dependent token groups are identified using strongly connected component algorithms, such as the Tarjan [tarjan] or Kosaraju-Sharir [sharir1981strong] algorithm.
There are two reasons for disregarding edges below the threshold. First, it helps to get rid of low dependencies often caused by noise. Second, since any two tokens within a single message are mentioned together at least once, a baseline of dependency always exists among all tokens of a message, leading to a fully connected interdependency graph. Therefore, since strongly connected components algorithms rely only on the presence of edges, not their weights, it will result in only one group.
@@ 3.6 Anchor Tokens Selection
After identifying strongly dependent token groups, the largest one is selected as the anchor group, and its members are designated as the anchor tokens. Algorithm 1 outlines all the steps required to select the anchor token.
@@ 3.7 Anchor-token-based Clustering
After identifying anchor tokens, messages are clustered into distinct groups based on their anchor tokens. Then, event IDs could be obtained from clusters by computing the cluster hash or its index. Should the desired outcome be solely the event IDs, the algorithm concludes at this juncture. However, if there is a request for templates or parameter masks in addition to event IDs, the algorithm proceeds to produce the next three stages.
@@ 3.8 Cluster Common Tokens
Common tokens within each cluster are those present across all messages in the cluster, effectively representing the intersection of tokens in all messages. Subsequently, these common tokens are later leveraged to generate event templates and parameter masks.
@@ 3.9 Event Templates
Tipping produces a set of templates for each event to handle different template variations of the event. For instance, in the example provided by Table I, all messages are representations of the same event. However, due to optional parameters within the log message structure[logpm], the templates are different from each other.
The process starts with a cluster of log messages and their common tokens. Initially, it involves processing these messages to create preliminary templates, where common tokens are kept and others are replaced with placeholders. Then, consecutive placeholders are merged into a single one. To make the template representable as a string, placeholders should be replaced by a parameter symbol, such as “<*>”. The final step aggregates these templates into a set, producing a set of templates for that cluster.
@@ 3.10 Parameter Masks
After identifying the common tokens, we initialize the parameter mask with an empty string for each message. Then, the parameter mask is populated by iterating over the message tokens, appending $n$ zeros if the token is present in the common tokens or $n$ ones otherwise, where $n$ is the number of characters in each token. Repeating this for all messages generates parameter masks for the entire cluster.
@@ 4 Experiments
Given that our algorithm relies on parallel computing, we implemented our algorithm in Rust due to its fearless concurrency, exceptional performance, and seamless interoperability with other languages. However, since Python is the dominant language in the data science and AI space, we also released a Python binding for our package (the core is still written in Rust), which will be utilized in the upcoming experiment. The Tipping Python package can be installed from pypi [^1] https://pypi.org/project/tipping, while the Rust package is available on crates.io [^2] https://crates.io/crates/tipping-rs.
Unless explicitly mentioned, all experiments are performed on a 64-bit OpenStack virtual machine featuring 16 Intel Broadwell 2 GHz virtual CPUs capable of boosting to high frequencies and 234 GiB of ECC memory, optimized for parallel processing and data integrity. All virtual disks are formatted utilizing EXT4 and XFS file systems for efficient data handling. Furthermore, the VM is equipped with two NVIDIA Tesla P100 GPUs (16 GB each). It is worth mentioning that none of the experiment methods utilizes a GPU.
We performed four experiments, see Sections 4.1 to 4.5, to answer the following research questions:
  • RQ1: How accurately and efficiently does Tipping perform in LogHub2k and LogHub2.0 benchmarks? See Section 4.1.
  • RQ2: How accurately and efficiently does Tipping perform in the LogPM benchmark? See Section 4.2.
  • RQ3: How sensitive is Tipping to its hyperparameters, such as $\theta$ and special patterns? See Section 4.3
  • RQ4: How fast is Tipping according to the LogLead benchmark? See Section 4.4
  • RQ5: What is the impact of Tipping as a parser on the downstream task? Experiment in Section 4.5
Finally, since our experiments yield multiple tables as results, yet they are presented in an aggregated format due to better readability and understandability, we have made the non-aggregated results available in a dedicated GitHub repository [^3] https://github.com/M3SOulu/tipping-detailed-result-tables for further transparency.
4.1 LogHub Benchmark
@@ 4.1.1 Experiment depiction
LogPai’s benchmark [loghubtools], published in 2019 and followed up by another study [loghub2] in 2024, has become a canonical accuracy evaluation benchmark for log parsers since its introduction. The first benchmark (LogHub2K) comprises 2,000 manually labeled samples across 16 public open datasets. It primarily uses Group Accuracy (GA) and the $F_{1}$ score as metrics. Additionally, it includes parsing time as a performance metric. However, it is observed that performance might not be fully representative of real-world scenarios in the first benchmark due to the limited number of samples in each dataset [logpm]. On the other hand, during the follow-up study [loghub2], the second benchmark (LogHub2.0) extended the datasets’ coverage. Furthermore, they added new metrics such as $F_{1}$ score of Group Accuracy (FGA) and Template Accuracy (TA), which contains Precision Template Accuracy (PTA), Recall Template Accuracy (RTA), and $F_{1}$ Template Accuracy (FTA).
In addition to the benchmark metrics, we introduced two new ones. The first metric, Parsed Dataset (PD), unveils the number of datasets that were successfully parsed without crashing or running out of memory within the time window defined by [loghub2] (12 hours).
The second metric, Normalized Parsing Time (NPT), reveals the total time spent on the benchmark (parsing all datasets), normalized by Tipping’s total time to demonstrate the parsing speed compared to Tipping.
In favor of transparency, we made the implementation, which is a fork of LogHub2.0, publicly available [^4] https://github.com/M3SOulu/loghub-2.0. Furthermore, we intend to merge this with the original repository upon acceptance.
@@ 4.1.2 Experiment results
In Table II, Tipping outperforms other methods across multiple key metrics. It achieves the highest Group Accuracy (GA) of 0.97 and $F_{1}$ score of Group Accuracy (FGA) of 0.93, far exceeding the averages of 0.71 and 0.58, respectively. Similarly, Tipping leads in Parsing Accuracy (PA) at 0.31, as well as in Precision Template Accuracy (PTA), Recall Token Accuracy (RTA), and $F_{1}$ score Template Accuracy (FTA).
TABLE II: The results of the LogHub 2k benchmark. The following acronyms are used: GA (Group Accuracy), FGA ($F_{1}$ score of Group Accuracy), PA (Parsing Accuracy), PTA (Precision Template Accuracy), RTA (Recall Template Accuracy), FTA ($F_{1}$ score of Template Accuracy), PD (Parsed Datasets), and NTP (Normalized Parsing Time).
Method GA FGA PA PTA RTA FTA PD NPT
AEL 0.81 0.68 0.28 0.22 0.26 0.23 14 1.19
Drain 0.85 0.74 0.31 0.28 0.29 0.28 14 3.78
IPLoM 0.77 0.77 0.20 0.17 0.16 0.16 14 2.75
LFA 0.65 0.65 0.22 0.14 0.14 0.13 14 2.53
LenMa 0.77 0.58 0.13 0.11 0.18 0.13 14 44.54
LogCluster 0.64 0.30 0.12 0.07 0.18 0.09 14 2.43
LogMine 0.73 0.45 0.21 0.13 0.18 0.13 14 25.77
LogSig 0.50 0.34 0.11 0.05 0.04 0.04 14 135.52
Logram 0.53 0.41 0.17 0.12 0.14 0.12 14 1.55
MoLFI 0.62 0.59 0.10 0.09 0.11 0.09 14 205.58
SHISO 0.68 0.65 0.09 0.13 0.13 0.13 14 19.12
SLCT 0.61 0.36 0.26 0.18 0.12 0.13 14 9.93
Spell 0.77 0.60 0.22 0.16 0.19 0.17 14 4.64
Tipping 0.97 0.93 0.31 0.36 0.35 0.35 14 1.00
Average 0.71 0.58 0.20 0.16 0.18 0.16 14 32.88
Furthermore, Tipping stands out in computational efficiency, achieving the fastest parsing time while other methods, on average, take 30 times more in the LogHub2k benchmark. This efficiency gap is even more widened when compared to MoLFI (205 times more) and LogSig (135 times more).
In Table III, Tipping continues to demonstrate strong performance, maintaining competitive metrics across the board in the LogHub2.0 benchmark. While its GA of 0.79 is lower than the top performer, AEL (0.86) performs better than the average (0.62). Its FGA is the highest at 0.74, being closely pursued by Drain (0.73), scoring significantly better than the 0.46 average. Furthermore, while not performing as the best, Tipping achieves a respectable over-average PA of 0.33. On the contrary, Tipping consistently shows top performance in the template accuracy metrics, including PTA (0.35), RTA (0.38), and FTA (0.36).
TABLE III: The results of the LogHub 2.0 benchmark. The following acronyms are used: GA (Group Accuracy), FGA ($F_{1}$ score of Group Accuracy), PA (Parsing Accuracy), PTA (Precision Template Accuracy), RTA (Recall Template Accuracy), FTA ($F_{1}$ score of Template Accuracy), PD (Parsed Datasets), and NPT (Normalized Parsing Time).
Method GA FGA PA PTA RTA FTA PD NPT
AEL 0.86 0.56 0.44 0.24 0.36 0.25 13 39.09
Drain 0.84 0.55 0.47 0.26 0.40 0.28 14 5.43
IPLoM 0.79 0.61 0.19 0.12 0.15 0.12 14 4.10
LFA 0.60 0.52 0.25 0.10 0.12 0.10 14 3.88
LenMa 0.82 0.47 0.24 0.14 0.26 0.16 10 120.13
LogCluster 0.57 0.04 0.11 0.01 0.16 0.01 14 2.05
LogMine 0.73 0.33 0.25 0.10 0.24 0.11 8 203.64
LogSig 0.18 0.16 0.09 0.01 0.01 0.01 11 98.71
Logram 0.34 0.20 0.20 0.04 0.24 0.05 10 136.49
MoLFI 0.59 0.35 0.11 0.03 0.09 0.04 11 106.59
SHISO 0.51 0.47 0.13 0.14 0.17 0.15 13 65.93
SLCT 0.40 0.27 0.24 0.10 0.23 0.11 11 87.88
Spell 0.70 0.33 0.22 0.07 0.16 0.08 12 102.86
Tipping 0.79 0.74 0.33 0.34 0.36 0.34 14 1.00
Average 0.62 0.40 0.23 0.12 0.21 0.13 12.07 69.84
In terms of parsing speed, Tipping shows the fastest performance. While being 69 times faster than the average, it is twice as fast as the second fastest (LogCluster). Furthermore, Tipping is among the five parsers that are capable of parsing all 14 datasets in the LogHub2.0 benchmark.
Overall, Tipping demonstrates promising results in both LogHub2k and LogHub2.0 benchmarks, producing comparatively accurate outputs with efficiency. This makes it a highly practical solution for both small-scale and large-scale log parsing.
@@ 4.2.1 Experiment depiction
LogPM benchmark [logpm] was introduced in response to the limitations of the LogHub benchmark. It distinguishes itself by emphasizing the detection of parameters at the character level. More in-depth, drawing inspiration from the field of computer vision, specifically semantic segmentation, LogPM introduces a new output for log parsers called the ”Parameter Mask.” This novel approach is accompanied by a new metric, the Parameter Mask Agreement (PMA), designed to compare the produced parsers’ parameter masks against the ground truth. Furthermore, LogPM endeavors to better replicate real-world prerequisites for performance evaluation by offering significantly larger datasets than those found in LogHub2k.
In favor of transparency, we merged our implementation of Tipping in the LogPM benchmark into the original publicly available LogPM repository [^5] https://github.com/M3SOulu/LogPMBenchmark.
@@ 4.2.2 Experiment results
In Table IV, Tipping demonstrates promising performance in the parsing accuracy metric in the LogPM benchmark. Tipping performs better in Android, Linux, and Proxifier datasets. In Zookeeper and HDFS, Tipping achieved the maximum accuracy alongside Lenma and Drain. Accordingly, Tipping performs either better or on par with the best performer in the competition. Furthermore, Tipping achieves the best average accuracy of 0.71, consolidating its place as the best overall performer in the PA metric.
TABLE IV: An in-depth comparison of Tipping’s performance with other leading state-of-the-art parsers in the LogPM benchmark, utilizing Parsing Accuracy as the evaluation metric.
@@
In Table V, presenting LogPM’s parameter mask agreement scores, Tipping emerges as a leading method with an average parameter mask agreement score of 0.87, showcasing its superior performance across various datasets. This performance is indicative of Tipping’s effectiveness in accurately identifying and matching parameter masks, a critical aspect of modern log parsing [logpmbench]. Comparatively, Brain and Lenma each hold an average agreement score of 0.80, indicating a commendable performance but still trailing behind Tipping. Both methods show strengths in certain areas, illustrating their potential in specific contexts. Drain and Spell, with average scores of 0.72 and 0.71, respectively, lag further behind, particularly struggling in datasets like Android.
TABLE V: An in-depth comparison of Tipping’s performance with other leading state-of-the-art parsers in the LogPM benchmark, utilizing Parameter Mask Agreement as the evaluation metric.
@@
In Table VI, Tipping performs the fastest in every dataset compared to its competition. The performance gap is more contrasted in larger datasets like HDFS, where Tipping performs 250 times better than the slowest parser (Drain). In terms of the averages, Tipping performs 5 times faster the the second best (Brain).
TABLE VI: An in-depth comparison of Tipping’s performance with other state-of-the-art parsers in the LogPM benchmark’s speed test. Other parsers’ times are divided by the Tipping time to demonstrate their speed compared to Tipping.
Dataset Brain Drain Lenma Spell Tipping
Android 2.58 25.89 5447.93 55.85 1.00
Apache 5.11 1.21 68.40 16.96 1.00
Hadoop 4.98 25.73 27.89 145.74 1.00
HDFS 8.55 257.82 34.18 147.77 1.00
HPC 3.13 1.58 15.14 41.88 1.00
Linux 4.75 4.36 45.58 116.52 1.00
Openstack 6.65 2.28 48.21 124.47 1.00
Proxifier 6.48 2.40 33.03 19.72 1.00
SSH 6.75 1.47 29.53 25.45 1.00
Zookeeper 5.62 2.13 27.81 25.77 1.00
Avg. 5.46 32.49 577.77 72.01 1.00
Overall, Tipping’s performance underscores its reliability and capacity to adapt to varied data types, making it a preferred choice for parsing tasks that are similar to LogPM.
4.3 Hyper-parameter Sensitivity
4.3.1 Experiment description
The previous experiments demonstrate Tipping’s performance across various benchmarks. However, its effectiveness could be questioned by the reliance on hyperparameters. In this experiment, we focus on investigating the impact of the threshold parameter, $\theta$, and the inclusion of special patterns (while and black).
To find the sensitivity of each hyperparameter, we repeat the LogHub2k and LogHub2.0 from Sec 4.1 and LogPM from Sec 4.2 with a variety of settings and analyze the results. More specifically, we repeat all three experiments with $\theta$ values in the range of $[0.1,1.0]$, both in the presence and absence of special tokens.
Figure 2 illustrates the sensitivity test results for LogHub2k. The results indicate that Tipping performs worse in the absence of regexes. Notably, the metrics exhibit consistent behavior in response to $\theta$ variations, regardless of whether regexes are used, with the exception of the extreme high-end. However, the performance gap attributed to regexes is more pronounced for GA than for TA.
4.3.2 Experiment results
Figure 2: LogHub2k sensitivity experiment for $\theta$ values in the range of $[0.1,1.0]$ and both with and without black and white special token regexes. The X-axis represents the $\theta$ value, while the Y-axis shows the respective metrics.
Similarly, Figure 3 showcases comparable findings for LogHub2.0. While the regex performance gap is less significant, the overall behavioral patterns remain consistent.
Figure 3: LogHub2.0 sensitivity experiment for $\theta$ values in the range of $[0.1,1.0]$ and both with and without black and white special token regexes. The X-axis represents the $\theta$ value, while the Y-axis shows the respective metrics.
Lastly, Figure 4 reveals similar observations for LogPM. Despite differences in metrics, the special tokens’ performance gap and the similarity in behavior across different values of $\theta$ follow the same pattern as the previous analysis.
Figure 4: LogPM sensitivity experiment for $\theta$ values in the range of $[0.1,1.0]$ and both with and without black and white special token regexes. The X-axis represents the $\theta$ value, while the Y-axis shows the respective metrics.
Overall, the observations bear witness to the fact that, except for the extreme ends, despite the possible gains from cherry-picking $\theta$ value and special tokens, the performance of Tipping is not entirely dependent on cherry-picked hyperparameters.
4.4 LogLead Computational Efficiency Benchmark